The mass and radius of a neutron star is 2MΘ and 15 km, respectively. Calculate the
value of gravitational red shift for light of wavelength 6000
oA
at a distance of 2 m
from its surface.
1
Expert's answer
2019-05-02T10:30:48-0400

Gravitational redshift is given by the following approximate equation


"\frac {\lambda} { \lambda_0}=1-\frac {GM} {r c^2} (1)"

where λ is the is the shifted wavelength, λ0 is the rest wavelength, G is the gravitational constant, and r is the distance from the gravitating body with mass M, c is the speed of light

Using (1) we got:


"\lambda=\lambda_0 (1-\frac {GM} {r c^2}) (2)"

In our case, 2MΘ=2×1.98892×1030 kg, r=2 m

We got: λ=881.4×10-6 m


Answer:

881.4×10-6 m